OpenStudy (anonymous):

How many total atoms are in 0.920 g of P2O5?

OpenStudy (anonymous):

First calculate moles of P2O5. moles = grams / molar mass The molar mass for P2O5 = 2 x P (2 x 31.0) + 5 x O (5 x 16.0) = 142 g/mole. 0.920 g P2O5 x (1 mole P2O5 / 142 g P2O5) = 0.0065 moles P2O5 0.0065 moles P2O5 x (6.02 x 10^23 molecules P2O5 / 1 mole P2O5) =3.913*10^23 molecules P2O5 Each molecule of P2O5 contains 2 P atoms and 5 O atoms (total = 7 atoms). 3.913 x 10^23 molecules P2O5 x (7 atoms / 1 molecule P2O5) =2.7391 x 10^26 atoms

OpenStudy (anonymous):

I don't think this is the right answer

OpenStudy (anonymous):

let me rework it

OpenStudy (anonymous):

I am in AP chemistry and this seems like basic chemistry problem to me

OpenStudy (anonymous):

It probably is I'm just awful at chemistry